OpenStudy (vishweshshrimali5):

Is 22051946 a perfect square ?

OpenStudy (vishweshshrimali5):

You cannot use calculator.

OpenStudy (anonymous):

it is not a perfect square

OpenStudy (vishweshshrimali5):

How did you get that ?

OpenStudy (anonymous):

22051946 = 2*11025973 so its odd * 2 cant be perfect

OpenStudy (vishweshshrimali5):

Great work guys. Thanks

OpenStudy (rational):

Clever :) how about this question @BSwan Is 22051946*2 a perfect square ?

OpenStudy (anonymous):

no :) but not the same reason hehe

OpenStudy (rational):

Is there an easy proof for that digital root method ?

OpenStudy (anonymous):

yep there is , mmm if u show n^2 is not of the form 3m or 3m+1 then its not perfect also there is a proof to know last digits for n^2 , wanna know it ?

mathslover (mathslover):

1 , 4, 6, 9 and 0 can be the last digits of n^2 ...!

OpenStudy (anonymous):

yes wanna prove it ?

mathslover (mathslover):

So, all of them are basically of the form 3m or 3m+1 3(0) + 1 = 1 3(1) + 1 = 4 3(2) = 6 3(3) = 9 3(0) = 0

OpenStudy (anonymous):

proof 1 :- 3k , 3k+1 proof 2 :- 0,1,4,5,6,9 they are independent

mathslover (mathslover):

What kind of proof are you referring to @BSwan ? Is it something logical, like n can have last digits as : 1 : n^2 = 1 2 : n^2 = 4 3 : n^2 = 9 4 : n^2 = 6 5 : n^2 = 5 6 : n^2 = 6 7 : n^2 = 9 8 : n^2 = 4 9 : n^2 = 1 0 : n^2 = 0 ( I just noticed that I missed 5 in the previous comment)

OpenStudy (anonymous):

number theory

mathslover (mathslover):

I have read the proof of it before, but probably forgot how it goes. Will love to see how you proove it. :-)

OpenStudy (anonymous):

okk :) let me get my notes , it will be better

mathslover (mathslover):

Sure. Take your time.

OpenStudy (anonymous):

so we start from here ok ? \(\large a=\sum_{n=0}^{k} a_n 10^{10}\)

mathslover (mathslover):

Okay.

OpenStudy (rational):

10^n *

OpenStudy (anonymous):

made a typo sry \(\large a=\sum_{n=0}^{k} a_n 10^{n}\)

OpenStudy (anonymous):

u know this cuz for any number a , digits a0,a1,a2....an cen be like this \(\large a=a_010^0+a_110^1+a_210^2+...+a_n10^n\) ok ?

OpenStudy (anonymous):

so take 10 common from terms -a_0 \( a=10 M +a_0 \) right ?

OpenStudy (rational):

\[\large a = a_ka_{k-1}\cdots a_1a_0 =\sum_{n=0}^{k} a_n 10^{n} = 10M + a_0\]

OpenStudy (anonymous):

take mod 10 for both sides \(\large a\mod 10 = 10 M \mod 10 + a_0 \mod 10\)

OpenStudy (rational):

that gives you last digit, so ?

OpenStudy (anonymous):

mmm rash where is 10 +

OpenStudy (anonymous):

yeah that gives u last digit for any number \(\large a = a_0 \mod 10\)

OpenStudy (anonymous):

nw , last digit for any number can be \(a_0\) =0,1,2,3,4,5,6,7,8,9

OpenStudy (rational):

\(\large a \equiv a_0 \mod 10\)

OpenStudy (anonymous):

nw to check n^2 \(\large a = a_0 \mod 10\) \(\large a^2 = a_0^2 \mod 10\) see ?

mathslover (mathslover):

Well, I need to learn modulo .. :(

OpenStudy (rational):

Yes ! last digit can only be {0, 1, 4, 5, 6, 9}

OpenStudy (rational):

how to prove the digital root stuff ?

OpenStudy (anonymous):

so we go again with \(a_0\) =1 up to 9 and check \(a_0\mod 10\) as u said before solver 0^2 mod 10 =0 . . . . 9^2 mod 10 =1 they would be the only sol

OpenStudy (rational):

Understood that last digit proof... I am trying to figure out how to prove restrictions on digital root

OpenStudy (rational):

@mathslover I can teach you mods in less than 30 minutes :)

OpenStudy (rational):

Will try some other time, ty @BSwan :)

OpenStudy (rational):

Oh wait actually it looks easy to prove with your summation notation

OpenStudy (rational):

nvm it looks complicated, wil try later >.<

mathslover (mathslover):

@rational thanks for that. May be, tomorrow or day after that, we will talk about this :-) Will make a post then we will discuss. Thanks again.

OpenStudy (anonymous):

sorry for laging ,i was having lunch :) lets proof digital root :)

ganeshie8 (ganeshie8):

Interesting, how are you going to express that process mathically ?

OpenStudy (anonymous):

to express lets say \(\S_K= \large \sum_{n=1}^{k} a_n \mod 10^n\) but im still thinking of a method to prove it